# Learning Neural Control Barrier Functions from Offline Data with Conservatism

### v1

@@ 4 Method
Our aim is to design a CBF-QP policy from offline data. If the reference policy or the dynamics are not available, we train corresponding models using the same data. Moreover, in settings with high-dimensional observations, we train representation models and consider the latent space to be the state space based on which we train the dynamics model and the neural CBF. We discuss how we train each of these components next.
4.1 Learning latent dynamics
@@
We train an autoencoder to project high-dimensional visual observations into a low-dimensional latent space. We train it in parallel with a control-affine latent dynamics model. The deterministic autoencoder maps RGB images $o_{t}\in\mathbb{R}^{c\times h\times w}$ to latent vectors $x_{t}\in\mathcal{X}$ using an encoder $\mathcal{E}$, and reconstructs them using a decoder $\mathcal{G}$. We train a control-affine dynamics model of the form:
@@
4.2 Learning nominal controllers
@@
We train neural reference controllers from offline data using behavior cloning (BC) and safe offline RL algorithms. We train BC-based controllers by solving the optimization problem $\operatorname*{argmax}_{\beta}\mathbb{E}_{(x,u)\sim\mathcal{D}}[\log\pi_{\beta}(u|x)]$. When we use only the safe trajectories in the dataset for behavior cloning, we call the resulting policy BC-Safe. For safety gymnasium environments [50], we train BC and safe offline‑RL policies using DSRL [51], which provides training datasets and reference implementations for safe offline‑RL algorithms.
4.3 Learning conservative control barrier functions from offline datasets
@@
We assume that a dataset of trajectories of the form $\sigma:=\{x_{1},a_{1},x_{2},a_{2},\dots,x_{T},a_{T}\}$ is available. We define $\mathcal{X}_{\text{safe}}$ as the set of safe states and $\mathcal{X}_{\text{unsafe}}$ as the set of unsafe states in the dataset. We denote the set of consecutive triplets $(x,a,x^{\prime})$ in the dataset in which $x^{\prime}\in\mathcal{X}_{\text{safe}}$ by $\mathcal{D}_{\text{safe}}$ and the set of triplets in which $x^{\prime}\in\mathcal{X}_{\text{unsafe}}$ by $\mathcal{D}_{\text{unsafe}}$. We also define $\mathcal{D}:=\mathcal{D}_{\text{safe}}\cup\mathcal{D}_{\text{unsafe}}$. Given this dataset, we train a neural CCBF $B_{\phi}:\mathcal{X}\to\mathbb{R}$ by minimizing the following loss function:
$\mathcal{L}_{\text{CCBF}}(\phi)$ $:=\mathcal{L}_{\text{safe}}(\phi)+\mathcal{L}_{\text{unsafe}}(\phi)+\mathcal{L}_{\text{ascent}}(\phi)+\mathcal{L}_{\text{descent}}(\phi)+\mathcal{L}_{\text{lip}}(\phi)+\mathcal{L}_{\text{c}}(\phi),$ (5)
where
$\mathcal{L}_{\text{safe}}(\phi)$ $:=\frac{w_{\text{safe}}}{|\mathcal{X}_{\text{safe}}|}\sum_{x\in\mathcal{X}_{\text{safe}}}\sigma(\varepsilon_{\text{safe}}-B_{\phi}(x)),\quad\mathcal{L}_{\text{unsafe}}(\phi):=\frac{w_{\text{unsafe}}}{|\mathcal{X}_{\text{unsafe}}|}\sum_{x\in\mathcal{X}_{\text{unsafe}}}\sigma(\varepsilon_{\text{unsafe}}+B_{\phi}(x)),$
@@
$\mathcal{L}_{\text{ascent}}(\phi)$ $:=\frac{w_{\text{ascent}}}{|\mathcal{D}_{\text{safe}}|}\sum_{(x,u,x^{\prime})\in\mathcal{D}_{\text{safe}}}\sigma(\varepsilon_{\text{ascent}}-\nabla B_{\phi}(x)\cdot(f_{\theta}(x)+g_{\theta}(x)u)-\gamma B_{\phi}(x)),$
@@
$\mathcal{L}_{\text{descent}}(\phi)$ $:=\frac{w_{\text{descent}}}{|\mathcal{D}_{\text{unsafe}}|}\sum_{(x,u,x^{\prime})\in\mathcal{D}_{\text{unsafe}}}\sigma(\varepsilon_{\text{descent}}+\nabla B_{\phi}(x)\cdot(f_{\theta}(x)+g_{\theta}(x)u_{\text{unsafe}})+\gamma B_{\phi}(x)),$
$\mathcal{L}_{\text{lip}}(\phi):=$ $\frac{w_{\text{lip}}}{|\mathcal{D}|}\sum_{(x,u,x^{\prime})\in\mathcal{D}}\left\|B_{\phi}(x^{\prime})-B_{\phi}(x)\right\|,\text{ and }\mathcal{L}_{\text{c}}(\phi):=\frac{w_{c}}{|\mathcal{X}_{\text{safe}}|}\sum_{x\in\mathcal{X}_{\text{safe}}}\left[\tau\ln\left(\sum_{v\in V_{x}}\exp\left(\frac{B_{\phi}(x^{\prime}_{v})}{\tau}\right)\right)\right].$
@@
Here, $\sigma$ is the ReLU activation function, and $\varepsilon_{\text{safe}}$, $\varepsilon_{\text{unsafe}}$, $\varepsilon_{\text{ascent}}$, $\varepsilon_{\text{descent}}$, $w_{\text{safe}}$, $w_{\text{unsafe}}$, $w_{\text{ascent}}$, $w_{\text{descent}}$, $w_{\text{lip}}$, $w_{c}$, and $\tau$ are hyperparameters. The set $V_{x}$ is a uniformly sampled set of actions from $\mathcal{U}$ to evaluate at state $x$. The state $x^{\prime}_{v}$ is the state reached by the learned dynamics following action $v$ at state $x$. The state $x^{\prime}$ is the next state following $x$ in the dataset, i.e., the one reached by system (1) following action $u$ at state $x$, where $u$ is the second element in the triple $(x,u,x^{\prime})\in\mathcal{D}$. When we train without using $\mathcal{L}_{\text{c}}$, we refer to the result as a neural control barrier function (NCBF).
The loss terms $\mathcal{L}_{\text{safe}}$ and $\mathcal{L}_{\text{unsafe}}$ drive the CCBF to be positive at safe states and negative at unsafe states. The terms $\mathcal{L}_{\text{ascent}}$ and $\mathcal{L}_{\text{descent}}$ drive the CCBF to satisfy the gradient constraint (2). Moreover, the loss term $\mathcal{L}_{\text{lip}}$ drives the learned CCBF to be Lipschitz continuous. Previous works on learning neural CBFs from data [52, 37, 11] have used only $\mathcal{L}_{\text{safe}}$, $\mathcal{L}_{\text{unsafe}}$ and $\mathcal{L}_{\text{ascent}}$ in their loss computation. We added the new term $\mathcal{L}_{\text{lip}}$ to smoothen the learned neural CBF. Finally, $L_{c}$ drives the CCBF to be conservative in its values on OOD states.
The term inside the sum in $\mathcal{L}_{\text{c}}$ represents the soft maximum of the CCBF values of the sampled states as well as the following state appearing in the training trajectory. The temperature parameter $\tau$ controls the smoothness of the soft maximum. As $\tau$ decreases, the soft maximum approaches the actual maximum. The gradient of $\mathcal{L}_{\text{c}}$ is equal to the gradient of $\frac{w_{c}}{|\mathcal{X}_{\text{safe}}|}\sum_{x\in\mathcal{X}_{\text{safe}}}\left[\tau\ln\left(\sum_{v\in V_{x}}\exp\left(\frac{B_{\phi}(x^{\prime}_{v})}{\tau}\right)\right)-B_{\phi}(x^{\prime})\right]$, but with the second term, $B_{\phi}(x^{\prime})$, detached from the computational graph during training in Pytorch using the tensor.detach() method. The latter is similar to the loss used in conservative Q learning (CQL) [23], where the Q-function is replaced by $B_{\phi}$, but CQL does not detach the second term. The reason we do so is to minimize $\mathcal{L}_{\text{c}}$ by decreasing the values in $\{B_{\phi}(x^{\prime}_{v})\}_{v\in V_{x}}$ instead of increasing the value of $B_{\phi}(x^{\prime})$. We want the CCBF values at InD states to be determined mainly by the loss terms $\mathcal{L}_{\text{safe}}$, $\mathcal{L}_{\text{unsafe}}$, $\mathcal{L}_{\text{ascent}}$, $\mathcal{L}_{\text{descent}}$, and $\mathcal{L}_{\text{lip}}$, which enforce the conditions required by the learned CBF to be a valid CBF, while minimally being affected by the conservative loss term $L_{c}$. We did an ablation study where we did not detach $B_{\phi}(x^{\prime})$. Its results, shown in Appendix E, support our argument that that the learned CCBFs without detaching $B_{\phi}(x^{\prime})$ generally under-perform those obtained by detaching it. Results of the latter are shown in Section 5.2.
5 Case Studies
@@
In this section, we describe the experimental setup across four environments and subsequently present our results along with the key observations.
5.1 Experimental setup
@@
We evaluate our approach on four environments: (1) 2D navigation: an agent must navigate to a goal position while avoiding a static circular obstacle, (2) Vision-based navigation: similar to 2D navigation but using bird-eye-view image observations, (3) SafetyHopper and (4) SafetySwimmer: locomotion tasks from the safety gymnasium benchmark, where agents must move as quickly as possible while adhering to velocity constraints. More details on these environments can be found in Appendix A. In each experiment, we present the results of rolling out the policy under different conditions: with no safety filters and with a NCBF-, an iDBF- and a CCBF-based ones. We compare these methods with FISOR [46] in the safety gymnasium tasks [50]. We train the NCBF, iDBF, and CCBF using the same hyperparameters including model size and only modify the hyperparameters related to the iDBF- and CCBF-specific loss terms.
@@
[…]
@@
This setup extends the 2D navigation task to a vision-based setting, where $64\times 64\times 3$ RGB observations representing a top-down view of the environment are given instead of the state, i.e., the 2D position.
@@
We test on SafetyHopper (11D state, 3D control) and SafetySwimmer (8D state, 2D control) from safety gymnasium [50]. Agents earn rewards for forward motion but incur costs if velocity exceeds user-defined thresholds. States are labeled unsafe if they exceed the maximum velocity constraint, and safe otherwise.
5.2 Results
We discuss two main observations: (1) CCBF-based filters significantly improves safety while minimally affecting performance, and (2) CCBF is easier to train and more robust to hyperparameters change than baselines. Further observations are discussed in Appendix C. We report the results for the 2D navigation and vision-based navigation scenarios in Table 1 using the metrics: success %, the percentage of episodes in which the agent successfully reached the goal; and collision %, the percentage of episodes with at least one collision. We visualize some of the NCBFs learned in the 2D navigation task in Figure 1 and provide additional visualizations in Figures 4(a) and 4(b) (in Appendix D). We report our results for the safety gymnasium tasks in Table 2 using the metrics: normalized reward and normalized cost, following the steps of previous works [51, 53, 54]. Detailed metric definitions can be found in Appendix A.4.2. Finally, Appendix B presents ablation studies investigating the effect of varying the density threshold $p$ for training the iDBFs as well as varying model sizes.

### v2

@@ IV Method
Our aim is to design a CBF-QP policy from offline data. If the reference policy or the dynamics are not available, we train corresponding models using the same data. Moreover, in settings with high-dimensional observations, we train representation models and consider their latent spaces to be the state spaces based on which we train the dynamics models and the neural CBFs. We discuss how we train each of these components next.
IV-A Learning latent dynamics
@@
We train a deterministic autoencoder to project high-dimensional visual observations into a low-dimensional latent space. We train it in parallel with a control-affine latent dynamics model. The autoencoder maps RGB images $o_{t}\in\mathbb{R}^{c\times h\times w}$ to latent vectors $x_{t}\in\mathcal{X}$ using an encoder $\mathcal{E}$, and reconstructs them using a decoder $\mathcal{G}$. We train a control-affine dynamics model of the form:
@@
Generally, we assume that the continuous time dynamics are either user-provided or that they can be approximated in the form of (5) accurately on the training distribution, which is the case in our experiments. Our assumption that the dynamics are control-affine is needed so that the online safety filtering problem can be formulated as a quadratic program.
IV-B Learning nominal controllers
@@
We train neural reference (or nominal) controllers on the same offline dataset we use for training the CCBF. We use behavior cloning (BC) and safe offline RL algorithms to train these controllers. We train BC-based controllers by solving the optimization problem $\arg\max_{\beta}\mathbb{E}_{(x,u)\sim\mathcal{D}}[\log\pi_{\beta}(u|x)]$. When we only use the safe trajectories in the dataset in BC, we call the resulting policy BC-Safe. In the navigation environment, we also use a PD controller defined as $\mathbf{u}_{\text{PD}}=K_{p}(\mathbf{x}_{g}-\mathbf{x})+K_{d}\frac{\mathbf{u}_{\text{nom}}}{\Delta t}$, where $\mathbf{x}_{g}$ is the goal position, and $K_{p},K_{d}$ are the proportional and derivative gains. For safety gymnasium environments [39], we train BC and safe offline RL policies using DSRL [40], which provides training datasets and reference implementations for safe offline RL algorithms. While such policies will be safer than the ones that result from traditional (non-safe) offline RL algorithms, they optimize for soft constraints and as we show in our experiments, generally become safer when coupled with a safety filter.
IV-C Learning CCBFs from offline datasets
@@
We assume that a dataset of trajectories of the form $\sigma:=\{x_{1},u_{1},x_{2},u_{2},\dots,x_{T},u_{T}\}$ is available. We define $\mathcal{X}_{\text{safe}}$ as the set of states labeled as safe and $\mathcal{X}_{\text{unsafe}}$ as the set of states labeled as unsafe in the dataset. We denote the set of consecutive triplets $(x,u,x^{\prime})$ in the dataset in which $x^{\prime}\in\mathcal{X}_{\text{safe}}$ by $\mathcal{D}_{\text{safe}}$ and the set of triplets in which $x\in\mathcal{X}_{\text{safe}}$ and $x^{\prime}\in\mathcal{X}_{\text{unsafe}}$ by $\mathcal{D}_{\text{unsafe}}$. Given this dataset, we train a neural CCBF $B_{\phi}:\mathcal{X}\to\mathbb{R}$ by minimizing the following loss function: $\mathcal{L}_{\text{CCBF}}(\phi)=\mathcal{L}_{\text{safe}}(\phi)+\mathcal{L}_{\text{unsafe}}(\phi)+\mathcal{L}_{\text{ascent}}(\phi)+\mathcal{L}_{\text{descent}}(\phi)+\mathcal{L}_{\text{lip}}(\phi)+\mathcal{L}_{\text{c}}(\phi)$, where $\mathcal{L}_{\text{safe}}(\phi)=\frac{w_{\text{safe}}}{|\mathcal{X}_{\text{safe}}|}\sum_{x\in\mathcal{X}_{\text{safe}}}\sigma(\varepsilon_{\text{safe}}-B_{\phi}(x))$,
$\mathcal{L}_{\text{unsafe}}(\phi)\!=\!\frac{w_{\text{unsafe}}}{|\mathcal{X}_{\text{unsafe}}|}\sum_{x\in\mathcal{X}_{\text{unsafe}}}\sigma(\varepsilon_{\text{unsafe}}+B_{\phi}(x))$,
@@
$\mathcal{L}_{\text{ascent}}(\phi)\!=\!\frac{w_{\text{ascent}}}{|\mathcal{D}_{\text{safe}}|}\sum_{(x,u,x^{\prime})\in\mathcal{D}_{\text{safe}}}\sigma(\varepsilon_{\text{ascent}}-\nabla B_{\phi}(x)(f_{\theta}(x)+g_{\theta}(x)u)-\alpha B_{\phi}(x))$,
@@
$\mathcal{L}_{\text{descent}}(\phi)=\frac{w_{\text{descent}}}{|\mathcal{D}_{\text{unsafe}}|}\sum_{\begin{subarray}{c}(x,u,x^{\prime})\in\mathcal{D}_{\text{unsafe}}\end{subarray}}\sigma(\varepsilon_{\text{descent}}+\nabla B_{\phi}(x)(f_{\theta}(x)+g_{\theta}(x)u)+\alpha B_{\phi}(x))$,
$\mathcal{L}_{\text{lip}}(\phi)\!=\!\frac{w_{\text{lip}}}{|\mathcal{D}|}\sum_{(x,u,x^{\prime})\in\mathcal{D}}\left\|B_{\phi}(x^{\prime})-B_{\phi}(x)\right\|$,
$\mathcal{L}_{\text{c}}(\phi)=\frac{w_{c}}{|\mathcal{X}_{\text{safe}}|}\sum_{x\in\mathcal{X}_{\text{safe}}}\tau\ln\left(\sum_{v\in V_{x}}\exp\left(\frac{B_{\phi}(x^{\prime}_{v})}{\tau}\right)\right)$.
@@
Here, $\sigma$ is the ReLU activation function, and $\varepsilon_{\text{safe}}$, $\varepsilon_{\text{unsafe}}$, $\varepsilon_{\text{ascent}}$, $\varepsilon_{\text{descent}}$, $w_{\text{safe}}$, $w_{\text{unsafe}}$, $w_{\text{ascent}}$, $w_{\text{descent}}$, $w_{\text{lip}}$, $w_{c}$, and $\tau$ are non-negative real numbers chosen as hyperparameters. The set $V_{x}$ is a uniformly sampled set of actions from $\mathcal{U}$ to evaluate at state $x$ as well as the action taken in the dataset at $x$. The state $x^{\prime}_{v}$ is the state reached by the learned dynamics following action $v$ at state $x$. When we train without using $\mathcal{L}_{\text{c}}$, we call the result a neural control barrier function (NCBF).
The loss terms $\mathcal{L}_{\text{safe}}$ and $\mathcal{L}_{\text{unsafe}}$ drive the CCBF to be positive at safe states and negative at unsafe states. The terms $\mathcal{L}_{\text{ascent}}$ and $\mathcal{L}_{\text{descent}}$ drive the CCBF to satisfy the gradient constraint (2). Specifically, $\mathcal{L}_{\text{descent}}$ penalizes violations of (2) during training. While it’s hard to determine violations of (2) inside the safe and unsafe sets, we can infer the violations whenever a trajectory crosses from a state labeled as safe to one that is labeled as unsafe. Moreover, the loss term $\mathcal{L}_{\text{lip}}$ drives the CCBF to be smooth. Finally, $\mathcal{L}_{c}$ drives the CCBF to be conservative.
The term inside the sum in $\mathcal{L}_{\text{c}}$ represents the soft maximum of the CCBF values of the sampled states as well as the next state appearing in the training trajectory. The temperature parameter $\tau$ controls the smoothness of the soft maximum. As $\tau$ decreases, the soft maximum approaches the actual maximum. When $\tau=1$, $\mathcal{L}_{\text{c}}$ is similar to a loss term appearing in the objective function of learning the $Q$-function for a learned policy $\pi_{k}$ in a variant of CQL described in equation (4) of [18]. That variant corresponds to choosing the sampling distribution $\mu$ in equation (4) in Section III-B of our paper to be the one that maximizes the argument of the argmin in the presence of a regularizer $\mathcal{R}(\mu)$, i.e.,
$\mu$ $\leftarrow\text{argmax}_{\mu}\alpha\,\mathbb{E}_{s\sim\mathcal{D},\,a\sim\mu(a|s)}[Q(s,a)]$
$\quad+\tfrac{1}{2}\,\mathbb{E}_{s,a\sim\mathcal{D}}\Big[Q(s,a)-\hat{B}^{\pi_{k}}\hat{Q}^{k}(s,a)\Big]^{2}+\mathcal{R}(\mu),$ (6)
where $\pi_{k}$ is the policy obtained by (IV-C) at iteration k1k-1, $\mathcal{R}(\mu)=-D_{\text{KL}}(\mu,\rho)$, $D_{\text{KL}}$ is the KL-divergence, and $\rho$ is the uniform distribution over the actions. Thus, $\mu$ would be the regularized maximizer of the iterate $\hat{Q}^{k}$. Substituting that $\mu$ in equation (4) in Section III-B results in:
$\hat{Q}^{k+1}$ $\;\leftarrow\;\arg\min_{Q}\;\alpha\,\mathbb{E}_{s\sim\mathcal{D}}[\log\sum_{a}\text{exp}(Q(s,a))]$
$\quad+\tfrac{1}{2}\,\mathbb{E}_{s,a\sim\mathcal{D}}\Big[Q(s,a)-\hat{B}^{\pi_{k}}\hat{Q}^{k}(s,a)\Big]^{2},$ (7)
as shown in equation (4) in [18]. In our setting, $\mathcal{L}_{\text{CCBF}}$ is analogous to the objective function in (7) where the $Q$-function is replaced with the barrier function $B_{\phi}$, the Bellman error term is replaced with the terms of $\mathcal{L}_{\text{CCBF}}$ besides $\mathcal{L}_{c}$ which penalize violations of the CBF conditions similar to how the Bellman error penalizes violations of the Bellman equation, and the sampling distribution for the actions in $V_{x}$ in $\mathcal{L}_{c}$ is chosen to be the policy that maximizes safety (the barrier values) with a regularizer in the form of the distance (KL-divergence) to the uniform distribution. Formalizing this connection is an interesting future direction.
V Case Studies
@@
In this section, we describe our experimental setup and present our results and key observations.
V-A Experimental setup
@@
We evaluate our approach on four environments: (1) 2D navigation: an agent must navigate to a goal position while avoiding a static circular obstacle, (2) Vision-based navigation: similar to 2D navigation but using bird-eye-view image observations, (3) SafetyHopper and (4) SafetySwimmer: locomotion tasks from the safety gymnasium benchmark [39], where agents must move as quickly as possible while adhering to velocity constraints. In each experiment, we present the results of rolling out the policy under different conditions: with no safety filter and with a NCBF-, an iDBF-, and a CCBF-based filter. We compare these methods with FISOR [33] in the safety gymnasium tasks. When training the NCBFs, iDBFs, and CCBFs, we use the same model architecture, the same set of common hyperparameters, and the same dynamics model. The only hyperparameters which differ between the different models are those specific to the different methods. Moreover, we train the NCBFs and CCBFs on the same datasets consisting of both safe and unsafe trajectories, whereas we train the iDBFs only on the safe states from safe trajectories in the dataset (also excluding the safe states in the unsafe trajectories), as we do for training BC-Safe policies.
@@
The agent has 2D single-integrator dynamics, has actuation limits that constrain the maximum speed in each dimension to be $v_{\mathit{max}}$, and has to reach a goal position while avoiding a static circular obstacle. Note that the agent can stop at any state by choosing zero control. Thus, the complement of the set of states that define the obstacle is a controlled forward invariant set. Consequently, we label a state as safe if the Euclidean distance $d$ to the center of the obstacle is greater than or equal to $r+\epsilon$, where $r$ is the obstacle’s radius and $\epsilon$ is a small positive constant. States with $d\leq r$ are labeled as unsafe, and states in between are not used for training. The small $\epsilon$ margin improves training stability, as has been observed in other works, e.g., [41].
@@
This setup extends the 2D navigation task to a vision-based setting, where $64\times 64\times 3$ RGB observations representing a top-down view of the environment are given instead of the 2D position.
@@
We test on SafetyHopper (11D state, 3D control) and SafetySwimmer (8D state, 2D control) from safety gymnasium [39]. Agents earn rewards for forward motion but incur costs if velocity exceeds user-defined thresholds. We picked these two tasks as challenging ones evident by the results in Table II which show that FISOR and other safe offline RL algorithms fail to balance the safety and task performance in them. We labeled states as unsafe if they exceeded the maximum velocity constraint, and labeled them as safe otherwise. Thus, we only labeled states in the failure set as unsafe and assumed that all other states in the dataset belong to a controlled forward invariant set. We also explored another strategy: in trajectories that eventually enter the failure set, we ignore the states outside the failure set and label only the states inside it (i.e., those with velocity exceeding the user-defined limit) as unsafe. States in trajectories that never enter the failure set over their entire duration are instead labeled as safe. However, this approach reduced the number of trajectories to sample safe states from to 135 out of 1686 trajectories for Swimmer and 130 out of 2240 for Hopper and led to poor safety-preserving performance. We summarize the key results of that experiment (detailed results in Table VII) and refer the readers to Section V-B for definitions of normalized cost (C) and reward (R): In Swimmer, averaged across all nominal controllers, using CCBF as a safety filter resulted in C=1.46 and R=0.31 (compared to C=1.49 and R=0.32 for nominal controllers without safety filter), outperforming NCBF (C: 2.38, R: 0.26). In the case of Hopper, both NCBF and CCBF performed poorly, degrading the performance of BC and BC-Safe, while maintaining the same performance for COptiDICE. BEAR-L and BCQ-L simulations terminated after only 5-30 steps because these algorithms, when paired with the CCBF and NCBF, produced policies that quickly drove the agent into terminal states such as falling, unstable torso angles, or extreme joint positions. Importantly, CCBF consistently outperformed NCBF under both labeling strategies, but both NCBF and CCBF achieved significantly better safety-preserving performance when trained using the full DSRL dataset compared to the trajectory-filtered approach for labeling safe states, emphasizing the importance of more data, even if it means potentially mislabeling few states.
V-B Results
We discuss two main observations: (1) CCBF-based filters significantly improve safety while minimally affecting performance, (2) CCBF is easier to train and more robust to hyperparameters changes than baselines. Then, we discuss the results o.ion studies that help us isolate the effect of our conservative loss term $\mathcal{L}_{c}$. We report the results for the 2D navigation and vision-based navigation scenarios in Table I using the metrics: success %, the percentage of episodes in which the agent successfully reached the goal; and collision %, the percentage of episodes with at least one collision. We report our results for the safety gymnasium tasks in Table II using the metrics: normalized reward and normalized cost, following the evaluation criteria in the safe offline RL literature [40, 42]. For both the Swimmer and the Hopper scenarios, they can can be modeled as CMDPs. We assume that the transition functions are deterministic and in the form of the control-affine dynamics in (5). Each episode results in a trajectory $\sigma:=\{x_{1},u_{1},r_{1},c_{1},\dots,x_{T},u_{T},r_{T},c_{T}\}$. The normalized reward and cost returns of a trajectory $\sigma$ generated using policy $\pi$ are defined as follows: $R_{\text{normalized}}^{\pi}:=\frac{R_{\sigma}^{\pi}-R_{\text{min}}}{R_{\text{max}}-R_{\text{min}}}\times 100$ and $C_{\text{normalized}}^{\pi}:=\frac{C_{\sigma}^{\pi}}{\kappa}$, where $R_{\sigma}^{\pi}=\sum_{t}r_{t}$ denotes the reward return of $\sigma$, $C_{\sigma}^{\pi}=\sum_{t}c_{t}$ denotes its cost return, and $R_{\text{min}}$ and $R_{\text{max}}$ denote the maximum and minimum reward returns of all the trajectories in the dataset as defined and computed in DSRL [40]. In all benchmark tasks, the cost $c_{t}$ is zero if the state $x_{t}$ is safe and one if it is unsafe. The Appendix contains the detailed tables of experimental results from our ablation studies, along with the hyperparameters used across all experiments.